# The Distribution of Atomic Hydrogen in the Host Galaxies of FRBs

### v2

@@ 4.3 Uncertainties
The uncertainties in the 1D metrics come from a number of sources; we first consider statistical contributions arising from the inherent noise component of each channel. We estimate this statistical uncertainty following Michalowski21; Glowacki_wallaby. We take the standard deviation of the line-free regions of the spectra at their final chosen resolutions, and remeasure each metric 1000 times with the spectra shifted by a noise array generated by a Gaussian of this width. The final uncertainties are taken as the standard deviation of these measurements.
@@
The 3DACs code does not output any uncertainties directly. Deg2023 claims that in their testing, the uncertainty associated with noise variation rarely exceeds 0.02, and almost certainly is dominated by systematics. We also note that the authors highlight imprecisions in the 3DACs results when the size of the Hi disk is smaller than four resolution elements (beams). Unfortunately, only three of our targets (FRB 20190425A, FRB 20200723B, and FRB 20211127I) satisfy this requirement. However, we still wish to investigate how useful these metrics are on real datasets across different instruments and resolutions. Thus, we choose to incorporate our own uncertainty associated with the choice of rotation centre — which is the quantity that introduces the most error due to low spatial resolution — by again retaking 1000 measurements with centres perturbed by a 3D Gaussian, with , , and all equal to 1. This is relatively straightforward as 3DACs already incorporates interpolation methods to allow for fractional rotation centres. While a Gaussian of this width likely overestimates the uncertainty, it is simple in design and clearly accounts for the effect of spatial and spectral resolution. Furthermore, as we do not present a statistical contribution, we are satisfied with the larger systematic estimate.
There are also uncertainties that arise due to the effects of galaxy inclination and viewing angle. While Deg2020 clearly shows that both have notable impacts on these measurements, we do not incorporate such considerations here as their contributions are nontrivial to model and likely require detailed kinematic modelling to accurately describe.
@@ 5 Target Analysis
Here we discuss the asymmetry results for each FRB host galaxy, highlighting interesting cases. Table 4 presents the quantified asymmetry metrics for each target; at the bottom, we include measurements for the known disturbed host of FRB 20171020A (Lee-waddell23) to compare against.
@@
FRB 20200723B | MeerKAT | 1.198 0.006 0.137 | 0.111 0.005 0.142 | 0.058 0.003 0.091 | 0.084 0.037 | 0.210 0.166 |
FRB 20210405I | MeerKAT | 1.135 0.060 0.069 | 0.300 0.043 0.076 | 0.170 0.026 0.042 | 0.074 0.066 | - |
FRB 20211127I | MeerKAT | 1.034 0.019 0.061 | 0.105 0.017 0.097 | 0.043 0.009 0.060 | 0.091 0.026 | 0.116 0.042 |
FRB 20211212A | MeerKAT | 1.244 0.125 0.299 | 0.371 0.086 0.143 | 0.140 0.048 0.065 | 0.126 0.055 | - |
@@
Targets with resolutions / SNRs great enough for 2D/3D measurements to be considered robust according to Deg2023.
@@ 5.1 FRB 20181220A
The host of FRB 20181220A is a highly inclined edge-on spiral which shows moderate Hi emission in both the GMRT and FAST data. By eye, the higher SNR FAST spectrum appears largely symmetric, which is reflected in its low asymmetry metrics. In contrast, the GMRT spectrum exhibits significantly greater asymmetry, though the deviation is consistent within the noise level. Furthermore, the intensity map shows a peak line flux away from the optical core of the galaxy. This causes the and 3D values to inflate to the highest in this sample. However, the area of the GMRT detection with respect to the beam size is right in the grey area of where Deg2023 identifies that 3D asymmetry modelling becomes less robust. Thus, we are hesitant to place too much weight on the GMRT result, especially considering the spectrum deviates from the FAST spectrum. The GMRT data does reveal that the galaxy lies in an Hi rich group, with at least two other nearby detections in the field, the nearest of which lies 0.55 Mpc away.
@@ 5.2 FRB 20181223C
This host is clearly significantly disturbed, though the origin of this morphology is difficult to ascertain. There are a number of galaxies in the visual field, including one very nearby to the south east, and another further away to the west. However, chimefrbs investigated these as potential hosts of the FRB, and found incompatible redshifts greater than 0.25 for both. Thus, there is no clear candidate that could be causing a disturbance through tidal means.
@@ 5.3 FRB 20190418A
With the second most distant redshift in this sample of 0.07132, the massive host of FRB 20190418A exhibits a weak Hi signal in the GMRT data. As such, its spectral features are largely dominated by the noise. However, the moment maps of Fig. 2 reveal an asymmetry in the extent of the galaxy either side of its core which gets picked up by SoFiA-2 even with a higher S+C threshold of 5. Furthermore, the asymmetry metrics all suggest disturbance; flux exceeds our threshold, spec is the highest in this sample, and even 1D — which is the most resilient to noise — is the also the highest in this sample. Thus, we consider this target disturbed, though deeper observations would be useful to confirm the nature of this.
@@ 5.5 FRB 20200723B
As the disturbance is spread across the entire northern face, and the velocity map shows no evidence of the more uniform bulk motion expected from a wind-driven outflow, this feature is most likely caused by a recent or ongoing minor tidal interaction. There are a number of nearby extended sources which may have gravitationally pulled on the outer gas, including the face-on spiral 2MASX J12404071-0505321, several fainter blobs that can be seen in the DECam imaging, and a few RC sources with no clear optical association. The tip of the outer contour in the 2532 km s-1channel map, which shows the most extreme elongation, appears to coincide with one of these extended sources. Unfortunately, with no further redshift information on the surrounding objects, it is not possible to definitively confirm this scenario. Regardless, these Hi observations reveal that NGC 4602 is another disturbed FRB host galaxy.
@@ 5.7 FRB 20211127I
The host of FRB 20211127I was previously analysed in Glowacki_wallaby using low resolution ASKAP data, and was prior to this study the only non-interacting FRB host observed in Hi. The detection appeared symmetric in spatial and spectral dimensions, but the authors concluded that higher resolution data would be required to fully confirm this case.
The new MeerKAT observation does indeed reveal a mostly symmetrical Hi distribution, both in the spectra and intensity map. The 1D asymmetry measurements are the lowest in this study, and while 3Dvalue returned by 3DACs is not -1, it is low. This is perhaps due to the slight elongation in the north east corner of the intensity map, but neither the channel nor velocity maps indicate anything definitive. Thus, we conclude that this host galaxy is undisturbed. Interestingly, it is worth noting that unlike the previous two undisturbed hosts, this galaxy is in a rich Hi group. Three nearby galaxies are detected in this deeper MeerKAT observation, none of which were present in the WALLABY Pilot Survey II observation of this field (Glowacki_wallaby). The nearest neighbour lies roughly 2.4 Mpc.
@@ 5.8 FRB 20211212A
The faint emission from this higher redshift host appears to be symmetrical, though neither the spectral nor spatial resolution is great enough to properly describe it. We conclude that it is mostly likely an undisturbed galaxy, but deeper observations would be needed to confirm this. This host has two faint Hi neighbours that each lie approximately 2 Mpc away.
@@ 5.11 FRB 20240201A
The neighbour is much brighter in Hi with its rotating arms and central bulge forming a strong triple peaked spectra. This dominates the FAST spectra, whose 3 arcminute beam easily encapsulates both whilst being centred on the smaller host. The slightly blueshifted nature of the Hi emission originating from the position of the FRB host, seen in the MeerKAT moment 1 map, confirms that it is not simply emission from the edges of the neighbour’s Hi cloud.
@@ 6.1 Are FRB Hosts Disturbed?
From this analysis, we find that six of these FRB host galaxies are disturbed: FRB 20181223C, FRB 20190418, FRB 20200723B, FRB 20231229A, FRB 20240201A, and FRB 20250316A. The hosts of FRB 20181220A, FRB 20201123A, and FRB 20231230A require deeper/targeted observations to draw conclusions, and the remaining four are undisturbed. With respect to the total population of FRB hosts now observed in Hi the confirmed ratio of disturbed to undisturbed galaxies is 11:4.
@@
Therefore, the observation that 70% of FRB hosts are disturbed cannot be directly contrasted to the 10-40% results in the literature. In Figure 9, we do utilise the flux and spec metrics to compare with the LVHIS and HALOGAS samples as is done in Michalowski21; Glowacki_hi_frb_whats_your_z, and also include approximations for the xGASS (xGASS) and HIPASS (Barnes2001) datasets. These are stellar mass and Hi mass selected samples that span a wide range of environments and masses, and thus are more representative of the local galaxy population. We estimate their flux values by extracting the data from Watts2020 and Reynolds2020 respectively. We note that robust comparison should be done exclusively with galaxies similar in mass and SFR to FRB hosts; however, both Watts2020; Reynolds2020LVHIS find no evolution in flux or spec as a function of mass, and thus comparing with these samples is satisfactory. Our sample of FRB hosts exhibits slightly higher 1D asymmetry levels than the background galaxy population, though the discrepancy is minimal — particularly when compared to the more representative background samples.
As discussed in Section 4.2, comparing our 1D, , and 3D values to literature is largely non-viable. Perron-Cormier2025 only includes figures of their 1D and 3D distributions for the 31% of galaxies in their sample which had 3D (i.e. not -1). Thus, these results are inherently biased against settled galaxies.
@@
To quantitatively assess the significance of this, we can assume Poisson error, and perform a likelihood ratio test. If our null hypothesis is that the ratio should follow 1:1 — i.e. the Poisson variable for both the disturbed and undisturbed host galaxy counts is 7.5 in this case — then a result of 11:4 results in a p-value associated with rejecting this hypothesis of 0.065. Compared to the state prior to this study (5:0 ratio gives a p-value of 0.008), the significance of the trend has decreased, with the p-value rising above the conventional 0.05 threshold. However, it remains on the cusp; in this low-number regime, a single count added to the disturbed and undisturbed sides would result in p-values of 0.041 and 0.129 respectively. As a result, a larger sample size of FRB hosts observed in Hi is required to clarify this tension.
Where does that leave the hypothesis regarding FRBs and the disruption in their hosts? It is tricky to say with certainty, though with four clear counterexamples, the connection between their progenitors and an increased star formation rate due to tidal disturbances is not consistent. This does not preclude the fast channel FRB progenitor theory, but, considering the borderline disturbance trend and the varied gas fractions of FRB hosts (see Sec. 6.2), it is unclear whether merger-induced star formation is a dominant driving factor in these galaxies as has been postulated in previous studies.
@@ 6.2 Hi Properties of FRB Hosts
As with most other galaxy population measurements, FRB hosts appear to span a broad range of values with respect to their global Hi properties. In Figure 10, we compare our hosts with the xGASS dataset (xGASS) and with previously analysed FRB hosts. Hosts of non-repeaters are shown in red, and hosts of repeaters are shown in blue. All of our new FRBs are non-repeaters, other than FRB 20200223B, whose host is our only non-detection. Measurements of SFR in the xGASS sample are derived from UV and mid-IR observations, tracing the past 100 Myrs of star-formation history (SFH). The values we use for our FRB hosts are taken from SED fitting in the respective localisation papers (see Table 1); the host of FRB 20210405I has not had SED fitting, and thus we use its RC-derived result, which should probe similar timescales of SFH. This host is shown with an unfilled star in panels (a) and (c).
Panel (a) shows that FRB hosts are mostly massive in Hi () with respect to the background population, which aligns with the trend seen in the optical literature (Gordon2023; Sharma2024). Interestingly, at face value there appears to be a slight separation between the repeater and non-repeater populations; the former all lie above the median line and are less massive in Hi whereas the latter mostly fall below the line and towards higher masses. Gordon2023; Sharma2024 both observed that repeater hosts may be less massive than non-repeater hosts, though due to the low number of repeaters in their samples (6 and 5), significance tests found no statistical evidence for this. However, Chen2025 presented a comparison between and SFR(H), and found a statistically significant separation (p-value of 0.0116) between the two populations when considering both properties in tandem.
Of the four repeating FRB hosts observed in Hi (the non-detection from this work and three from archival data), only one is included in those studies. We run a simple two sample Kolmogorov-Smirnov (KS) test on the Hi masses of our sample, and also run a MANOVA test (as in Chen2025) on the multivariate domain of HI vs. SFR. In each case, we find -values greater than 0.05 (0.27 and 0.17 respectively), though this is largely expected due to the small sample size. Regardless, this now indicates yet another example of by-eye discrepancy in the masses of repeaters and non-repeaters; future Hi studies of FRB hosts should consider probing this relation a focus.
@@
Panel (c) displays a more robust relationship between gas fraction and specific SFR (sSFR), used by Kaur2022; Glowacki_wallaby to assess if any large gas mass had recently been deposited into the galaxy through a merger. FRB hosts tend to lie above the median for Hi detected xGASS galaxies, but the spread again is quite large. The highly interacting host of FRB 20180916B (upper blue diamond) lies further above its respective bin median than any in our new sample; however, the two disturbed hosts presented in Michalowski21 fall below this median, demonstrating that this relationship is not necessarily a consistent indicator of overall disturbance.

### v3

@@ 4.3 Uncertainties
The uncertainties in the 1D metrics come from a number of sources; we first consider statistical contributions arising from the inherent noise component of each channel. We estimate this statistical uncertainty following Michalowski21 (66, 31). We take the standard deviation of the line-free regions of the spectra at their final chosen resolutions, and remeasure each metric 1000 times with the spectra shifted by a noise array generated by a Gaussian of this width. The final uncertainties are taken as the standard deviation of these measurements.
@@
The 3DACs code does not output any uncertainties directly. Deg2023 (24) claims that in their testing, the uncertainty associated with noise variation rarely exceeds 0.02, and almost certainly is dominated by systematics. We also note that the authors highlight imprecisions in the 3DACs results when the size of the Hi disk is smaller than four resolution elements (beams). Unfortunately, only three of our targets (FRB 20190425A, FRB 20200723B, and FRB 20211127I) satisfy this requirement. However, we still wish to investigate how useful these metrics are on real datasets across different instruments and resolutions. Thus, we choose to incorporate our own uncertainty associated with the choice of rotation centre — which is the quantity that introduces the most error due to low spatial resolution — by again retaking 1000 measurements with centres perturbed by a 3D Gaussian, with , , and all equal to 1. This is relatively straightforward as 3DACs already incorporates interpolation methods to allow for fractional rotation centres. While a Gaussian of this width likely overestimates the uncertainty, it is simple in design and clearly accounts for the effect of spatial and spectral resolution. Furthermore, as we do not present a statistical contribution, we are satisfied with the larger systematic estimate.
There are also uncertainties that arise due to the effects of galaxy inclination and viewing angle. While Deg2020 (23) clearly shows that both have notable impacts on these measurements, we do not incorporate such considerations here as their contributions are nontrivial to model and likely require detailed kinematic modelling to accurately describe.
@@ 5 Target Analysis
Here we discuss the asymmetry results for each FRB host galaxy, highlighting interesting cases. Table 4 presents the quantified asymmetry metrics for each target; at the bottom, we include measurements for the known disturbed host of FRB 20171020A Lee-waddell23 (52) to compare against.
@@
| FRB 20200723B | MeerKAT | 1.198 0.058 0.137 | 0.111 0.037 0.142 | 0.058 0.023 0.091 | 0.084 0.037 | 0.210 0.166 |
| FRB 20210405I | MeerKAT | 1.135 0.148 0.069 | 0.300 0.099 0.076 | 0.170 0.026 0.042 | 0.074 0.066 | - |
| FRB 20211127I | MeerKAT | 1.034 0.060 0.061 | 0.105 0.056 0.097 | 0.043 0.037 0.060 | 0.091 0.026 | 0.116 0.042 |
| FRB 20211212A | MeerKAT | 1.244 0.563 0.299 | 0.371 0.222 0.143 | 0.140 0.048 0.065 | 0.126 0.055 | - |
@@
Targets with resolutions / SNRs great enough for 2D/3D measurements to be considered robust according to Deg2023 (24).
@@ 5.1 FRB 20181220A
The host of FRB 20181220A is a highly inclined edge-on spiral which shows moderate Hi emission in both the GMRT and FAST data. By eye, the higher SNR FAST spectrum appears largely symmetric, which is reflected in its low asymmetry metrics. In contrast, the GMRT spectrum exhibits significantly greater asymmetry, though the deviation is consistent within the noise level. Furthermore, the intensity map shows a peak line flux away from the optical core of the galaxy. This causes the and 3D values to inflate to the highest in this sample. However, the area of the GMRT detection with respect to the beam size is right in the grey area of where Deg2023 (24) identifies that 3D asymmetry modelling becomes less robust. Thus, we are hesitant to place too much weight on the GMRT result, especially considering the spectrum deviates from the FAST spectrum. The GMRT data does reveal that the galaxy lies in an Hi rich group, with at least two other nearby detections in the field, the nearest of which lies 0.55 Mpc away.
@@ 5.2 FRB 20181223C
This host is clearly significantly disturbed, though the origin of this morphology is difficult to ascertain. There are a number of galaxies in the visual field, including one very nearby to the south east, and another further away to the west. However, chimefrbs (9) investigated these as potential hosts of the FRB, and found incompatible redshifts greater than 0.25 for both. Thus, there is no clear candidate that could be causing a disturbance through tidal means.
@@ 5.3 FRB 20190418A
With the third most distant redshift in this sample, the massive host of FRB 20190418A exhibits a weak Hi signal in the GMRT data. As such, its spectral features are largely dominated by the noise. However, the moment maps of Fig. 3 reveal an asymmetry in the extent of the galaxy either side of its core which gets picked up by SoFiA-2 even with a higher S+C threshold of 5. Furthermore, the asymmetry metrics all suggest disturbance; flux exceeds our threshold, spec is the highest in this sample, and even 1D — which is the most resilient to noise — is also the highest in this sample. Thus, we consider this target disturbed, though deeper observations would be useful to confirm the nature of this.
@@ 5.5 FRB 20200723B
As the disturbance is spread across the entire northern face, and the velocity map shows no evidence of the more uniform bulk motion expected from a wind-driven outflow, this feature is most likely caused by a recent or ongoing minor tidal interaction. There are a number of nearby extended sources in the DECam imaging, though the only Hi detection visible in MeerKAT arises from a large spiral 1.6 Mpc to the south. The tip of the outer contour in the 2532 km s-1channel map, which shows the most extreme elongation, appears to overlap with a faint extended source, though this is likely a coincidence. Unfortunately, with no further redshift information on the surrounding objects, it is not possible to definitively characterise the system. Regardless, these Hi observations reveal that NGC 4602 is another disturbed FRB host galaxy.
@@ 5.7 FRB 20211127I
The host of FRB 20211127I was previously analysed in Glowacki_wallaby (31) using low resolution ASKAP data, and was prior to this study the only non-interacting FRB host observed in Hi. The detection appeared symmetric in spatial and spectral dimensions, but the authors concluded that higher resolution data would be required to fully confirm this case.
The new MeerKAT observation does indeed reveal a mostly symmetrical Hi distribution, both in the spectra and intensity map. The 1D asymmetry measurements are the lowest in this study, and while 3Dvalue returned by 3DACs is not -1, it is low. This is perhaps due to the slight elongation in the north east corner of the intensity map, but neither the channel nor velocity maps indicate anything definitive. Thus, we conclude that this host galaxy is undisturbed. Interestingly, it is worth noting that unlike the previous two undisturbed hosts, this galaxy is in a rich Hi group. Three nearby galaxies are detected in this deeper MeerKAT observation, none of which were present in the WALLABY Pilot Survey II observation of this field Glowacki_wallaby (31). The nearest neighbour lies roughly 2.4 Mpc away.
@@ 5.8 FRB 20211212A
The faint emission from this high redshift host appears to be symmetrical, though neither the spectral nor spatial resolution is great enough to properly describe it. We conclude that it is mostly likely an undisturbed galaxy, but deeper observations would be needed to confirm this. This host has two faint Hi neighbours that each lie approximately 2 Mpc away.
@@ 5.11 FRB 20240201A
The spectra presented in Fig. 2 show the superposition of both galaxies’ emission. The host contributes partially to the first peak but is dominated by the neighbour; examining the resolved MeerKAT data reveals the neighbour’s rotating arms form the bulk of the outer peaks and its notably gas rich central bulge forms the third peak in the centre. The slightly blueshifted nature of the Hi emission originating from the position of the FRB host, seen in the MeerKAT moment 1 map, confirms that it is not simply emission from the edges of the neighbour’s Hi cloud.
@@
5.12 FRB 20240210A
The spiral host of FRB 20240210A lies only 1.5’ from of a smaller neighbour to the south west, and both exhibit clear Hi emission. When imaging with a Briggs weighting with a robust value of 2, SoFiA-2 considers the galaxies to be connected. However, the significance of any potential tidal bridge is extremely marginal; even when rebinned from a spectral resolution of 26 kHz to 104 and 208 kHz resolutions, the link never rises above 2, and is completely absent when imaging with a robust weighting of 0.5. Furthermore, individual channel maps do not suggest any consistent extension of the host galaxy.
Given the neighbour’s archival redshift of 0.02360 from the 2dF Galaxy Redshift Survey 2dFGRS (20), the angular separation deprojects into a physical distance of roughly 1 Mpc. The spectral profile of the host is symmetric, and none of the asymmetry metrics hint at intrinsic disruption. As such, we do not consider this host to be tidally interacting with the neighbour, and thus is undisturbed.
5.13 FRB 20240312D
The host of FRB 20240312D shows an unusual gas distribution in its intensity map, with the emission extending further to the south than to the north. Channel maps, presented in Fig. 11, confirm this asymmetry, as the southern extension is present across at least a third of the velocity profile. However, these features are less significant than in previous hosts: they do not reach the high contour levels seen in the host of FRB 20181223C, nor are they highly resolved by the beam as in the host of FRB 20200723B, making the evidence less conclusive. This is further reflected in the host’s 3D value of 0.199, which sits above every undisturbed host and below every disturbed host.
This galaxy resides in a highly rich Hi group containing at least eight nearby Hi galaxies, a number of which are within 1 Mpc. Cross-matching these detections with DECam imaging reveals that, apart from one case, they are all small, faint dwarf galaxies. The larger edge-on spiral visible to the south in Fig. 11 appears close in projection, but the central frequency of its strong Hi line places it nearly 10 Mpc in the background. Consequently, the nearest companion to the FRB host lies roughly 600 kpc to the east. Given the host exists in such a dense environment, and the observed gas extension does not point toward the nearest neighbour, it is more likely that the asymmetry is simply a lingering imprint of an earlier gravitational perturbation caused during the formation of the group than a recent tidal interaction. In these cases, peripheral gas can take considerable time to resettle into the disk. We therefore regard this host as a middle-ground case and, in Section 6, examine it under both classifications when interpreting the results.
5.14 FRB 20240615B
The Hi emission arising from this distant host appears largely symmetrical. The line is narrow, but the double-horned profile is still resolved, with the peaks reaching similar intensities. Neither its moment maps nor its asymmetry metrics suggest any disturbance, and thus we conclude this host is settled. It is worth noting that this galaxy does lie in a relatively Hi rich group, with at least four galaxies within 10 Mpc, the nearest positioned roughly 900 kpc away.
@@ 6.1 Are FRB Hosts Disturbed?
From this analysis, we find that six of these FRB host galaxies are disturbed: FRB 20181223C, FRB 20190418A, FRB 20200723B, FRB 20231229A, FRB 20240201A, and FRB 20250316A. Another six are clearly undisturbed: FRB 20190425A, FRB 20210405I, FRB 20211127I, FRB 20211212A, FRB 202404210A, and FRB 20240615B. One is borderline (FRB 20240312D), and the remaining three require deeper/targeted observations to draw firm conclusions. With respect to the total population of FRB hosts now observed in Hi the confirmed ratio of disturbed to undisturbed galaxies is 11:6.
@@
Therefore, the observation that 65% of FRB hosts are disturbed cannot be directly contrasted to the 10-40% results in the literature. In Figure 12, we do utilise the flux and spec metrics to compare with the LVHIS and HALOGAS samples as is done in Michalowski21 (66, 30), and also include approximations for the xGASS xGASS (12) and HIPASS Barnes2001 (6) datasets. These are stellar mass and Hi mass selected samples that span a wide range of environments and masses, and thus are more representative of the local galaxy population. We estimate their flux values by extracting the data from (Watts2020, 88, Fig. 2 ) and (Reynolds2020, 74, Table 5 ) respectively. We note that robust comparison should be done exclusively with galaxies similar in mass and SFR to FRB hosts; however, both Watts2020 (88, 75) find no evolution in flux or spec as a function of mass, and thus comparing with these samples is satisfactory. Our sample of FRB hosts exhibits slightly higher 1D asymmetry levels than the background galaxy population, though the discrepancy is minimal — particularly when compared to the more representative background samples.
As discussed in Section 4.2, comparing our 1D, , and 3D values to literature is largely non-viable. Perron-Cormier2025 (69) only includes figures of their 1D and 3D distributions for the 31% of galaxies in their sample which had 3D (i.e. not -1). Thus, these results are inherently biased against settled galaxies.
@@
To quantitatively assess the significance of this, we can assume Poisson error, and perform a likelihood ratio test. If our null hypothesis is that the ratio should follow 1:1 — i.e. the Poisson variable for both the disturbed and undisturbed host galaxy counts is 8.5 in this case — then a result of 11:6 results in a p-value associated with rejecting this hypothesis of 0.222. This is far higher than the standard significance threshold of 0.05; even considering the borderline case of FRB 20240312D as a disturbed host only shifts the p-value to 0.153. As such, we find that the earlier apparent trend suggesting underlying disturbance in FRB host galaxies is no longer statistically significant. While our observations, which reveal that 21 of 22 FRB hosts emit in Hi, further link FRB progenitors with star formation, the data provide no evidence that merger-induced star formation enhancement plays a significant role in their production.
@@ 6.2 Hi Properties of FRB Hosts
As with most other galaxy population measurements, FRB hosts appear to span a broad range of values with respect to their global Hi properties. In Figure 13, we compare our hosts with the xGASS dataset xGASS (12) and with previously analysed FRB hosts. Hosts of non-repeaters are shown in red, and hosts of repeaters are shown in blue. All of our new FRBs are non-repeaters, other than FRB 20200223B, whose host is our only non-detection. Measurements of stellar mass and SFR in the xGASS sample are derived from UV and mid-IR observations, tracing the past 100 Myrs of star-formation history (SFH). The values we use for our FRB hosts are taken from SED fitting in the respective localisation papers (see Table 1). The hosts of FRB 20210405I, 20240312D, and 20240615B have not had SED fitting, and thus we derive stellar mass approximations using their 2MASS -band magnitudes and the mass-luminosity relation for spiral galaxies McGaugh2014 (61), and use their RC-derived SFRs, which should probe similar timescales of SFH to SED fitting; these hosts are shown with unfilled stars in Fig. 13.
Panel (a) shows that FRB hosts have standard SFRs for their Hi masses, but are mostly massive in Hi () with respect to the background population, which aligns with the stellar mass trend observed in optical studies Gordon2023 (34, 81). Interestingly, the four repeater hosts appear to lie towards the less massive end of the FRB host Hi mass range; Gordon2023 (34, 81) also noted a similar pattern in their stellar masses. However, due to the low number of repeaters in their samples (6 and 5), significance tests found no statistical evidence for such a trend. However, Chen2025 (15) presented a comparison between and SFR(H) in 11 repeater and 33 non-repeater hosts, and found a statistically significant separation (p-value of 0.0203 when excluding high redshift non-repeaters) between the two populations when considering both properties in tandem.
Only one of the repeater hosts in our sample was included in any of the stellar mass studies above, potentially providing further evidence to separate repeater and non-repeater host populations based on mass. However, with such a small number of targets, and given the level of overlap observed in panel (a), this phenomenon is clearly not yet significant. To verify this, we run a simple two sample Kolmogorov-Smirnov (KS) test on the Hi masses of our sample, and also run a MANOVA (Multivariate Analysis of Variance) test (Chen2025, 15, as in) on the multivariate domain of HI vs. SFR. In each case, we find -values greater than 0.05 (0.11 and 0.16 respectively). Regardless, future Hi studies of FRB hosts should consider probing this relation a focus.
@@
Panel (c) displays a more robust relationship between gas fraction and specific SFR (sSFR), used by Kaur2022 (47, 31) to assess if any large gas mass had recently been deposited into the galaxy through a merger. FRB hosts tend to lie above the median for Hi detected xGASS galaxies, but the spread again is quite large. The highly interacting host of FRB 20180916B (upper blue diamond) lies further above its respective bin median than any in our new sample; however, the two disturbed hosts presented in Michalowski21 (66) fall below this median, demonstrating that this relationship is not necessarily a consistent indicator of overall disturbance.